OpenStudy (anonymous):

2x5 - 9x4 + 12x3 - 12x2 + 10x - 3 = 0 whats the answer?

OpenStudy (anonymous):

\[2 x^5-9 x^4+12 x^3-12 x^2+10 x-3=(x-3) (x-1) (2 x-1) \left(x^2+1\right) \]